Subject: 51641 6apr00 SDSS night log Date: Fri, 7 Apr 2000 06:51:00 -0600 (MDT) From: sjnk@apo.nmsu.edu (Scot Kleinman) APACHE POINT OBSERVATORY SDSS 2.5M OBSERVING LOG Thursday April 6 2000 (MJD 51641) OBSERVING TEAM: =============== Dan Long 1st shift (APO) Stephanie Snedden 1st shift (APO) Scot Kleinman 2nd shift (APO) Atsuko Kleinman 2nd shift (APO) Eric Neilsen PT (JHU) Visitors: THE PLAN: ========= Image stripe 37 N until about lambda of 15. Then image the Groth strip ~1 hour each of 32 N and 32 S, lambda of 10 to 30. Spectroscopy is set to go if conditions are not suitable for imaging. SYNOPSIS: ========= Stripe 37 N imaged for about 1.5 hours from survey starting lambda Data quality questionable for 37 N due to FFS petals getting in the way. Imaging interrupted to tighten flat field petals. Imaging 37 N resumed after about an hour, and continued for another 1.5 hours in marginal seeing. Switched to spectroscopy at 01:00. We got 4 exposures (done) on plate 301 and 3 exposures on plate 314. endNight ran successfully- anyone know why? GENERAL COMMENTS (see Problems below for details): ================================================== We started out the night with the imager mounted. During twilight we focused and set up. At astronomical twilight we started a science scan at the beginning of stripe 37 N. We may have a problem with findRotator. After processing the first lskip, findRotator will come up with a number. Implementing the rotator change and testing that number with a second lskip, findRotator comes up with a number which is roughly equal and opposite to that from the first requested lskip. However, issuing a third lskip without moving the rotator, findRotator shows the rotator to be right on. So does a fourth lskip/findRotator. This is the second night in a row when the setup for the first scan of the night did this. We were very careful tonight to wait for rotator moves to clear and grab the correct frames, but the same thing happened. So last night was apparently not operator error ( PR 968 ) [Second shift editorial: we got email about this recently that this has been discussed by others as well and there is a significantly larger variation in the findRotator reported value than the errors it gives suggests. Multiple lskips are the only way to go. Sigh.] Tonight tccmon caused a telrun slowdown. The logging feature was enabled and when tccmon was exited a whole slew of TCC broadcast data scrolled by and telrun resumed its normal alacrity. We noticed several occurrences of sharp dips in sky brightness which looked a lot like clouds do on the scrolling display, yet the cloud camera showed it to be perfectly clear. Suspecting the wind was causing problems we went down to the telescope. After a couple of minutes a gust came by and slammed a couple of the flat field petals closed. After the gust died down, it re-opened. We ended the drift and slewed the telescope to stow. After a call to French to confirm how to adjust the clutches, we tightened them all down. It is now significantly harder to open or close them by hand. We did not see a repeat of them flapping in the breeze despite gusts of 25 MPH. ( PR 969 ) 37 N was resumed after about an hour, and scanned for about an hour and a half. Seeing stubbornly remained above 1.5 arcseconds, so when the next available spectro cartridge was about an hour from prime integration time, the decision was made to end the drift and switch to spectroscopy. This was at about 01:00. The instrument change took just under an hour thanks in large part to the awkwardness of loading the corrector onto a cartridge. There was one cartridge already on the cart, in the front position, so we couldn't get a new cartridge next to it and there's no room in the garage to put it away. So we put the corrector on this cartridge (which is time consuming in and of itself -especially in the dark) and mounted it, then withdrew to the manipulator to move the cartridge to the back bay, then put the cartridge we wanted on the cart and mounted it. We couldn't have moved the unwanted cartridge to the other cart bay first since the corrector only goes in one position with two cartridges on the cart. There may have been a more efficient way to move things around, but the point is we shouldn't have to worry about such things. The present situation is very awkward and time consuming. My (Scot's) suggestion is to have a dummy cartridge and its own cart (and shed) where we keep the corrector- this way it never gets lifted by anything but the cartridge underneath it and we don't have to worry about finding a space for it in the garage or manipulating it at all! We (I've talked to at least some of the other observers about this) like this idea better than the planned cathouse mini-manipulator thing that is planned. Its own cart and own mounting cartridge means it never gets picked up by anything or anybody except for the telescope lift. Make sense, yes? For cart 8 (and 6), we took two sets of starPattern data- one right before the science exposures, and the latter, after the diamondPoints. The results are listed in the spectro data section. The last one (maybe two) exposure (maybe exposures) occurred after the "optimal" viewing time as shown in the plug plate database. Can someone please give us feedback on the quality of these exposure(s) so we can decide, in general, if it's worth continuing past optimal to save the setup and calibration time, or if we really need to eat it and do the exposures again along with the new overhead? During the end of night dewar-a-rama, we broke an air hose connector that sits right by the base of the spectro dewar. We know this has happened at least once before. French was on his way in and said he'd look at it then. We ran endNight from SOP (3_12_2) and it finished without errors! Why is this? Is sop's endNight different from iop's? If so, shall we run endNight from SOP the rest of this run until we test and debug the new fixes in the latest version of iop? Tapes ---- Node Ship to FNAL Stay at APO id1 2510 2516 id2 2511 2517 id3 2512 2518 id4 2513 2519 id5 2514 2520 id6 2515 2521 Gang 2522 2523 DATA ACQUISITION (see IOP/SOP output below): =========================================== IMAGING: -------- 1361 Bias ~20 frames 1362 setup for 37 S 1363 Science 37 S startLambda -53.5 at about 21:00 raBoreErr: -12.7 +- 2.7", decBoreErr: -8.0 +- 1.7" frames 20 29: azErr: -19.4 +- 4.3", elErr: -10.6 +- 2.3" muErr: -12.0 +- 1.4", nuErr: 5.4 +- 2.4" Ended around 22:30, due to trouble with the flatfield screens not staying open in the brisk winds. 1364 setup on 37 again after ffs adjustment. 1365 Science 37 S startLambda -17 at about 23:40. raBoreErr: -16.4 +- 4.7", decBoreErr: 4.7 +- 1.7" Frames 50 to 60 azErr: -4.9 +- 3.2", elErr: -7.9 +- 1.8" muErr: -4.3 +- 1.9", nuErr: 7.2 +- 1.8" Significant drop in sky brightness in r ( and some in g ) seen from frames 60 to 120. The az went from about 195 to 185 with alt of 55. The Hollomann effect? Stopped 1365 at lambda of 3 at 01:00. Seeing did not improve and the decision was made to switch to spectroscopy. SPECTROSCOPY: (sensor cables unplugged) --------------------------------------- Cartridge 8 plate 301 --------------------- Frame flavor ..... ...... 3317 -bias -init 3318 flat 3319 arc gFiber *rot* gFiber(gprobe,1,rot) = -720 gFiber(gprobe,10,rot) = -9.5704620274724999 gFiber(gprobe,11,rot) = -80.50733566871051 gFiber(gprobe,2,rot) = -169.75056516752804 gFiber(gprobe,3,rot) = -45.00078053990083 gFiber(gprobe,4,rot) = 171.25910475332969 gFiber(gprobe,5,rot) = 13.331381814160455 gFiber(gprobe,6,rot) = 46.452575718384729 gFiber(gprobe,7,rot) = 18.121761282035632 gFiber(gprobe,8,rot) = -56.779315893967791 gFiber(gprobe,9,rot) = 161.3271709927626 3320 science 900s 3321 science 900s 3322 science 900s 3323 science 900s 3324-3327 diamondPoints gFiber(gprobe,1,rot) = 158.15961249209138 gFiber(gprobe,10,rot) = -14.312889475940203 gFiber(gprobe,11,rot) = -84.046024107477365 gFiber(gprobe,2,rot) = -152.68687364426577 gFiber(gprobe,3,rot) = -48.443940782256526 gFiber(gprobe,4,rot) = 176.71900305461008 gFiber(gprobe,5,rot) = 2.8118992315408273 gFiber(gprobe,6,rot) = 32.495431471392202 gFiber(gprobe,7,rot) = 13.007816820511476 gFiber(gprobe,8,rot) = -59.734196010439419 gFiber(gprobe,9,rot) = 155.12507386991942 3328 flat 3329 arc cartridge 6 plate 314 --------------------- exposure type -------- ---- 3330 pre-preCalib flat (telescope not in position: used to get guider flat) gFiber(gprobe,1,rot) = 77.405205053559783 gFiber(gprobe,10,rot) = 70.414522362627238 gFiber(gprobe,11,rot) = -79.584661852629637 gFiber(gprobe,2,rot) = 91.821009065322215 gFiber(gprobe,3,rot) = 110.44974150260458 gFiber(gprobe,4,rot) = 77.445128825551507 gFiber(gprobe,5,rot) = 177.22485235668117 gFiber(gprobe,6,rot) = -164.87259158645085 gFiber(gprobe,7,rot) = -116.11196018202045 gFiber(gprobe,8,rot) = -104.50955036033575 gFiber(gprobe,9,rot) = -90.920535552087543 3331 flat 3332 arc 3333 science 900s 3334 science 900s 3335 science 900s - this plate extended a few minutes into twilight 3336-3339 diamondPoints gFiber(gprobe,1,rot) = 76.200409658745755 gFiber(gprobe,10,rot) = 70.575450883300675 gFiber(gprobe,11,rot) = -79.24897828359974 gFiber(gprobe,2,rot) = 92.091604176761763 gFiber(gprobe,3,rot) = 108.7551867456539 gFiber(gprobe,4,rot) = 73.401752510476697 gFiber(gprobe,5,rot) = 177.91145024373029 gFiber(gprobe,6,rot) = -162.77995081847791 gFiber(gprobe,7,rot) = -116.63377757407815 gFiber(gprobe,8,rot) = -108.55492549791548 gFiber(gprobe,9,rot) = -89.662020499647767 3340 flat 3341 arc MIRROR NUMBERS ============== *** Info taken at nominal primary piston *** Primary: setmir( cmd ): 0.00 -3.9 3.9 213.5 213.5 galil( act ): 4850 -2350 100 -9500 50 -100 linear gauges Tonight Nominal 1 = Axial A = 1.9315 1.9315 2 = Axial B = 0.5570 0.5570 3 = Axial C = 0.5570 0.5570 4 = Transverse = 0.4090 0.4075 5 = East Lateral = 1.8985 1.8980 6 = West Lateral = 0.3985 0.3990 Secondary: ( inst=imager focus=-144 alt= 65.3 deg T=9.0C) setmir( act ): -416.77, 5.01, -5.03, 50.16, -113.08 galil ( act ): -512893., -525844., -536137., 2000., -2800. linear gauges 1 = Axial A = 0.349 2 = Axial B = 0.386 3 = Axial C = 0.376 4 = Transverse = 0.071 FOCUS LOG: ========== M1 Piston M2 Piston Temp Wind Time Inst setmir TCC setmir Az Alt C Dir MPH filt FWHM " 21:30 Imager 0 -144 -417 256 65.3 9.0 279 18 23-i' 1.41 22:00 Imager 0 -136 -404 236 62.9 9.0 285 21 23-i' 1.45 00:00 Imager 0 -163 -407 197 54.5 7.7 314 21 23-i' 1.78 00:45 Imager 0 -166 -410 187 54.7 7.9 305 17 23-i' 1.62 02:50 Spectro 0 -100 1055 -24 54.8 8.0 329 21 1.8 05:40 Spectro 0 -100 1086 -50 45.1 7.3 13 25 1.8 WEATHER SYNOPSIS ================ Tonight was clear with west to north winds. WEATHER DATA: ============= Wind Time Temp F Dewp F MPH Direction Dust DIMM Sky 16:01 55 11 20 285 (WNW) 615 - Clear 17:01 56 12 11 264 (W) 682 - Clear 18:01 56 13 15 291 (WNW) 716 - Clear 19:01 53 11 11 270 (W) 733 - Clear 20:01 50 9 14 262 (W) 832 - Clear 21:01 49 7 20 266 (W) 775 - Clear 22:01 48 6 17 271 (W) 708 - Clear 23:01 47 8 22 293 (WNW) 960 - Clear 00:01 46 10 20 313 (NW) 1034 - Clear 01:01 47 11 5 335 (NNW) 1165 - Clear 02:01 46 9 12 351 (N) 777 1.0 Clear 03:01 46 12 14 325 (NW) 835 1.0 Clear 04:01 45 13 9 16 (NNE) 875 1.0 Clear 05:01 45 13 12 6 (N) 849 1.0 Clear LN2 SCHEDULE: ============= Imager dewar changed at 06:00 7Apr00. Spectrographs filled at 19:30 and 06:00 and dewar replaced. TELESCOPE STATUS: ================= Spectro cartridge 6 plate 314 mounted (and corrector). INTERLOCKS: ----------- None bypassed. SOFTWARE USED: -------------- IOP/SOP: v3_12_2 Watcher: v2_8_2 MCP: v4_2_1 TPM: v1_0_7 AstroDa: 14_20 PROBLEMS IN DETAIL: =================== Yowza! DATA LOGS FROM IOP/SOP: =======================